OpenStudy (anonymous):

write an equation for the line that is parallel to the given line and passes through the given point. y=x+8; (2,6)

OpenStudy (anonymous):

y-6=(x-2)

OpenStudy (anonymous):

choices are y=x+6 or y=x-6

OpenStudy (anonymous):

well then, the coordinate i know may give the answer y=x+4 :P

OpenStudy (anonymous):

??